OpenStudy (anonymous):

Solve the Equation Show all teps 1) a/4 + 2 =-6 2)2ns/5 + 4 =0 3)5+x/6=-6 the slash means its a fraction

OpenStudy (softballgirl372015):

a + 2 = -6(4) a + 2 = -24 a = -24 - 2 a = -26

OpenStudy (anonymous):

how about 2 and 3